# EventFlow: Forecasting Continuous-Time Event Data with Flow Matching

### v1

@@ 5.2 Unconditional Generation of Event Sequences
Next, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution. This task serves as a benchmark to evaluate the methods in terms of how well they are able to fit the underlying TPP. Moreover, learning a general-purpose TPP prior could enable downstream tasks, such as data augmentation (Graikos et al., 2022).
In Tables 2 and 3, we report MMD values (11) for each of the synthetic and real-world datasets. We tune the hyperparameters for all models and perform model selection based on the validation set MMD. MMDs are calculated by sampling 1,000 sequences from each trained model, and estimating Equation (11) using the generated and test set samples. The first row (“data”) is the MMD calculated between samples in the training and validation sets, giving us a sense of the best-case performance.
Overall, we find that our EventFlow method (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets. This is particularly pronounced on the real-world datasets, where we obtain the lowest MMD on 5 of the 7 datasets. We see that IFTPP (mean rank: $\mathbf{2.1}$) is a strong baseline, obtaining results which are competitive with our method. The Add-and-Thin method (mean rank: $\mathbf{2.4}$) is often similarly strong, but struggles on the SC dataset. While the NHP (mean rank: $\mathbf{3.7}$) can obtain good fits, this appears to be dataset dependent, with weak results on the NSR, SC, and Reddit-S datasets. The diffusion baseline (mean rank: $\mathbf{4.8}$) is our weakest baseline, which is perhaps unsurprising as this model can only be trained to maximize the likelihood of a subsequent event and not the overall sequence likelihood.
Hawkes1 Hawkes2 NSP NSR SC SR
Data $1.3$ $1.3$ $1.8$ $3.0$ $5.7$ $1.1$
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
Data $1.3$ $0.6$ $0.4$ $3.1$ $2.6$ $3.6$ $3.1$
@@ 6 Conclusion
In this work, we propose EventFlow, a non-autoregressive and likelihood-free generative model for temporal point processes. We demonstrate that EventFlow is able to achieve state-of-the-art results on a multi-step forecasting task and strong performance on unconditional generation.
There are several directions in which our work could be extended. First, we do not explicitly enforce that the support of our model TPP is $[0,T]$. This would necessitate moving beyond the Gaussian setup, which is non-trivial to carry out. Second, more sophisticated approaches to learning the event count distribution $p_{\phi}(n\mid\mathcal{H})$ could lead to improved performance. Overall, we believe that our work lays a foundation for future flow-based TPPs, and exploring their application in tasks such as imputation, marked TPPs, and querying (Boyd et al., 2023) are exciting directions.
Reproducibility Statement
Our experiments were conducted with fixed random seeds and tested to ensure reproducibility. After publication, we will open-source our code to facilitate reproducibility.
@@ Appendix A Datasets
In this section, we provide some additional details regarding the datasets used in this work. In Table 4, we report the number of sequences in each dataset, some basic statistics regarding the number of events in each sequence, and their support $[0,T]$ and chosen forecast window $\Delta T$. In all datasets, we use 60% of the data for training, 20% for validation, and the remaining 20% for testing.
@@ Appendix C EventFlow Architecture and Training Details
Here, we provide additional details regarding the parametrization and training of our EventFlow model. In general, our model is based on the transformer architecture (Vaswani, 2017; Yang et al., 2022), due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions. In all settings, our reference measure $\mu_{0}$ is specified with $q=\mathcal{N}(0,I)$.
@@
For our unconditional model, we first embed the sequence times $\gamma_{s}$, the flow-time $s$, and the sequence position indices using sinusoidal embeddings followed by an additional linear layer. There are three linear layers in total – one for the flow time, one shared across the sequence times, and one for the position indices. These embeddings are added together to create a representation of the sequence, and we apply a standard transformer to this sequence to produce a sequence of vectors of length $N(\gamma_{s})$. Finally, each of these vectors is projected to one dimension via a final linear layer with shared weights to produce the vector field $v_{\theta}(\gamma_{s},s)$. See Figure 2.
@@
Our architecture for predicting the number of future events given a history, i.e. $p(n\mid\mathcal{H})$, is again based on the transformer decoder, sharing the same overall architecture as our unconditional model. However, the key difference is that we instead take a mean of the final sequence embeddings before passing this through a small MLP to produce the final logit. See Figure 4.
@@
We normalize all sequences to the range $[-1,1]$, using the overall min/max event time seen in the training data. All sequences are generated on this normalized scale, prior to re-scaling the sequence back to the original data range before evaluation. Our model is trained with the Adam (Kingma & Ba, 2015) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler (Loshchilov & Hutter, 2017), which cycled every $10,000$ steps. Final hyperparameters were selected by best performance on the validation dataset achieved at any point during the training, where models were evaluated 10 times throughout their training.
@@
Figure 4: Overview of our architecture modeling the event count distribution $p_{\phi}(n\mid\mathcal{H})$. The model takes as input an observed history $\mathcal{H}$. As in our other architectures, the events are embedded and passed through a transformer. Here, the final sequence embedding output by the transformer is averaged and passed through an additional residual MLP with three layers to produce the logit corresponding to $p(n\mid\mathcal{H})$.
@@
Table 7: The best hyperparameter settings found for the event count predictor $p(n\mid\mathcal{H})$ in our EventFlow method on the forecasting task.
Learning Rate Emb. Dim. MLP Dim. Heads Transformer Layers
PUBG $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
Reddit-S $10^{-3}$ $128$ $256$ $4$ $6$
Taxi $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Twitter $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Yelp-Airport $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Yelp-Miss. $5\times 10^{-4}$ $512$ $2048$ $8$ $6$

### v2

@@ 5.2 Unconditional Generation of Event Sequences
Finally, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution. In Table 3 we report MMD values for each of the synthetic and real-world datasets. See Appendix E for results with standard deviations. EventFlow (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets. While IFTPP (mean rank: $\mathbf{2.1}$) performs well on the synthetic datasets, it struggles with the more complex real-world datasets, reflecting its high bias. The Add-and-Thin method (mean rank: $\mathbf{2.4}$) is often similarly strong, but struggles on the SC dataset. While the NHP (mean rank: $\mathbf{3.7}$) can obtain good fits, this appears to be dataset dependent, with weak results on the NSR, SC, and Reddit-S datasets. The diffusion baseline (mean rank: $\mathbf{4.8}$) is our weakest baseline, which is perhaps unsurprising as this model can only be trained to maximize the likelihood of a subsequent event and not the overall sequence likelihood.
Table 3: MMDs (1e-2) between the test set and $1,000$ generated unconditional sequences averaged over five random seeds. EventFlow achieves the overall highest mean rank.
H1 H2 NSP NSR SC SR PUBG Red.C Red.S Taxi Twitter YelpA YelpM
IFTPP $\mathbf{1.5}$ $\mathbf{1.4}$ $\mathbf{2.3}$ $\underline{6.2}$ $\mathbf{5.8}$ $\mathbf{1.3}$ $5.7$ $1.3$ $1.9$ $5.8$ $\mathbf{2.9}$ $8.2$ $\underline{5.1}$
NHP $\underline{1.9}$ $5.2$ $3.6$ $12.6$ $25.4$ $5.0$ $7.2$ $2.2$ $22.5$ $\underline{5.0}$ $7.3$ $6.7$ $6.1$
Diffusion $4.8$ $5.5$ $10.8$ $15.0$ $9.1$ $5.1$ $14.3$ $3.9$ $6.2$ $11.7$ $12.5$ $10.9$ $10.5$
Add-Thin $\underline{1.9}$ $2.5$ $\underline{2.6}$ $7.4$ $22.5$ $2.2$ $\underline{2.8}$ $\underline{1.2}$ $2.7$ $5.2$ $\underline{4.8}$ $\mathbf{4.5}$ $\mathbf{3.0}$
EventFlow (25 NFEs) $\underline{1.9}$ $\underline{2.2}$ $3.8$ $\mathbf{4.2}$ $\underline{8.3}$ $\underline{1.7}$ $\mathbf{1.5}$ $\mathbf{0.7}$ $\mathbf{0.7}$ $\mathbf{3.5}$ $4.9$ $\underline{6.6}$ $\mathbf{3.0}$
@@ 6 Conclusion
In this work, we propose EventFlow, a non-autoregressive generative model for temporal point processes. We demonstrate that EventFlow is able to achieve state-of-the-art results on a multi-step forecasting task and strong performance on unconditional generation. Our work lays a foundation for future flow-based TPPs, and exploring applications in tasks like imputation, marked TPPs, and querying (Boyd et al., 2023) are exciting directions.
Limitations and Outlook
In this work, we have focused on modeling the event times, and thus do not experiment with marked TPPs or spatiotemporal TPPs. Extending EventFlow to these tasks is a promising directions. We also note that we do not explicitly enforce that the support of our model is $[0,T]$, which would necessitate moving beyond the Gaussian setting. Moreover, our baseline methods use the neural architectures proposed in the original works, which vary across the methods. Carefully studying the effects of these architecture choices is an important direction.
@@ Appendix A Datasets
In this section, we provide some additional details regarding the datasets used in this work. In Table 4, we report the number of sequences in each dataset, some basic statistics regarding the number of events in each sequence, and their support $[0,T]$ and chosen forecast window $\Delta T$. In all datasets, we use 60% of the sequences for training, 20% for validation, and the remaining 20% for testing.
@@
For selecting the forecasting horizon $\Delta T$, we follow the choice made in Lüdke et al. (2023) to facilitate comparison and reproducibility. We note the initial times $T_{0}\in[\Delta T,T-\Delta T]$ are selected to be at least $\Delta T$ to avoid forecasting sequences with no observations in the history.
@@ Appendix B Proofs
In this section, we provide a proof of Proposition 1, showing a necessary and sufficient condition for the existence of balanced couplings.
@@ Appendix C EventFlow Architecture and Training Details
Here, we provide additional details regarding the parametrization and training of our EventFlow model. In general, our model is based on the transformer architecture (Vaswani et al., 2017; Yang et al., 2022), due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions. We emphasize that this is an effective architecture choice, but our method is not necessarily tied to this architecture, i.e., other choices like CNNs or RNNs are certainly possible here. In all settings, our reference measure $\mu_{0}$ is specified with $q=\mathcal{N}(0,I)$. All models (including the baselines) were trained on a cluster of six NVIDIA A6000 GPUs with 24Gb of VRAM. No hardware parallelization was used, i.e., each model was only trained on a single GPU.
@@
Our unconditional model takes in a sequence $\gamma_{s}=(t^{1}_{s},\dots,t^{n}_{s})$ and flow time $s$. We first embed the sequence times $t^{k}$, the flow-time $s$, and the sequence position indices. These position indices are handled by sinusoidal embeddings followed by an additional linear layer. There are three linear layers in total – one for the flow time, one shared across the sequence times, and one for the position indices. These embeddings are added together to create a representation for each element of the sequence, and we apply a standard transformer to this sequence to produce a sequence of vectors of length $N(\gamma_{s})$. Finally, each of these vectors is projected to one dimension via a final linear layer with shared weights to produce the vector field $v_{\theta}(\gamma_{s},s)$. See Figure 3.
@@
Our architecture for predicting the number of future events given a history, i.e. $p_{\phi}(n\mid\mathcal{H})$, is again based on the transformer decoder, sharing the same overall architecture as our unconditional model. However, the key difference is that we instead take a mean of the final sequence embeddings before passing this through a small MLP to produce the final logit. See Figure 5. We note that this requires specifying a maximum value $N_{\max}$, which we set as the maximum sequence length seen in the training data. In early experiments, we also parametrized the model not through a logit but rather as a mixture of Poissons or a mixture of Negative Binomials. While this no longer has the limitation of assuming a value for $N_{\max}$, we found that this approach yielded worse results.
@@
We normalize all sequences to the range $[-1,1]$, using the overall min/max event time seen in the training data. All sequences are generated on this normalized scale, prior to re-scaling the sequence back to the original data range before evaluation. Our models are trained with the Adam (Kingma & Ba, 2015) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler (Loshchilov & Hutter, 2017), which cycled every $10,000$ steps. Final hyperparameters were selected by best performance on the validation dataset achieved at any point during the training, where models were evaluated 10 times throughout their training.
@@
To train the event count model $p_{\phi}(n\mid\mathcal{H})$, we seek to minimize the regularized cross-entropy loss
$\mathcal{L}(\phi)=\mathbb{E}_{n,\mathcal{H}}\left[-\log p_{\phi}(n\mid\mathcal{H})+\frac{\alpha}{N_{\max}}\sum_{k=1}^{N_{\max}}p_{\phi}(k\mid\mathcal{H})(n-k)^{2}\right]$ (19)
where $\mathcal{H}$ is a given history and $n$ is the number of events in $\mathcal{T}$ following this history. Note that this is a cross-entropy loss where the regularization term can be viewed as a squared optimal transport distance between the distribution $p_{\phi}(n\mid\mathcal{H})$ and a delta distribution $\delta[n]$ at the true $n$. This regularization term encourages the values of $p_{\phi}(n\mid\mathcal{H})$ to be smooth in $n$, as it penalizes predictions which are far fronm the true $n$ more heavily than those which are nearby. We searched for values of $\alpha$, the weight associated with the OT-loss, over the set $\{0,1/N_{\max},10/N_{\max},100/N_{\max},1000/N_{\max}\}$.
@@
Figure 5: Overview of our architecture modeling the event count distribution $p_{\phi}(n\mid\mathcal{H})$. The model takes as input an observed history $\mathcal{H}$. As in our other architectures, the events are embedded and passed through a transformer. Here, the final sequence embedding output by the transformer is averaged and passed through an additional residual MLP with three layers to produce the logit corresponding to $p_{\phi}(n\mid\mathcal{H})$.
@@
Table 7: The best hyperparameter settings found for the event count predictor $p_{\phi}(n\mid\mathcal{H})$ in our EventFlow method on the forecasting task.
Learning Rate $\alpha\times N_{max}$ Emb. Dim. MLP Dim. Heads Transformer Layers
PUBG $10^{-3}$ $1.0$ $512$ $2048$ $8$ $6$
Reddit-C $10^{-3}$ $1000.0$ $128$ $256$ $4$ $6$
Reddit-S $10^{-3}$ $1000.0$ $128$ $256$ $4$ $6$
Taxi $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
Twitter $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
Yelp-Airport $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
Yelp-Miss. $10^{-3}$ $100.0$ $512$ $2048$ $8$ $6$
@@ Appendix D Additional Details on Baselines
Appendix E Additional Experiments
This section contains additional empirical evaluations, supplementing Section 5.
E.1 MARE
In Table 16, we evaluate the performance of the various models only in terms of the predicted number of events in the forecast. To do so, we measure the mean absolute relative error (MARE) given by
$\mathrm{MARE}=\mathbb{E}_{n,\mathcal{H},\hat{n}\sim p_{\phi}(n\mid\mathcal{H})}\left|\frac{\hat{n}-n}{n}\right|$ (20)
where $n$ represents the true number of points in a sequence following a history $\mathcal{H}$ and $\hat{n}\sim p_{\phi}(n\mid\mathcal{H})$ represents the predicted number of points. The expectation is estimated empirically on the testing set. As our method directly predicts the number of events $n$ by sampling from the learned distribution $p_{\phi}(n\mid\mathcal{H})$, this serves as a direct evaluation of this component of our model. We note that the baselines only model $n$ implicitly. For example, the autoregressive models sample new events until an event is generated outside the support $\mathcal{T}$. As an ablation, we also report the MARE values for our method without regularization, i.e., setting $\alpha=0$. We see that the regularization significantly improves the event count predictions.
Table 16: MARE values evaluating the predicted number of events when forecasting. Mean values $\pm$ one standard deviation are reported over five random seeds. The lowest MARE on each dataset is indicated and bold, and the second lowest is indicated by an underline.
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
IFTPP $1.05{\scriptstyle\pm 0.14}$ $1.69{\scriptstyle\pm 0.39}$ $0.79{\scriptstyle\pm 0.20}$ $0.60{\scriptstyle\pm 0.11}$ $0.88{\scriptstyle\pm 0.08}$ $0.76{\scriptstyle\pm 0.07}$ $0.76{\scriptstyle\pm 0.05}$
NHP $1.02{\scriptstyle\pm 0.08}$ $\underline{0.95}{\scriptstyle\pm 0.01}$ $1.00{\scriptstyle\pm 0.0004}$ $0.67{\scriptstyle\pm 0.11}$ $2.48{\scriptstyle\pm 0.40}$ $0.80{\scriptstyle\pm 0.22}$ $1.07{\scriptstyle\pm 0.34}$
Diffusion $1.95{\scriptstyle\pm 0.48}$ $1.28{\scriptstyle\pm 0.09}$ $1.12{\scriptstyle\pm 0.56}$ $0.49{\scriptstyle\pm 0.07}$ $0.66{\scriptstyle\pm 0.04}$ $0.65{\scriptstyle\pm 0.07}$ $0.72{\scriptstyle\pm 0.07}$
Add-and-Thin $\underline{0.43}{\scriptstyle\pm 0.01}$ $0.99{\scriptstyle\pm 0.10}$ $\underline{0.38}{\scriptstyle\pm 0.01}$ $\underline{0.33}{\scriptstyle\pm 0.02}$ $\underline{0.60}{\scriptstyle\pm 0.02}$ $\mathbf{0.42}{\scriptstyle\pm 0.01}$ $\mathbf{0.46}{\scriptstyle\pm 0.03}$
EventFlow (ours) $\mathbf{0.40}{\scriptstyle\pm 0.01}$ $\mathbf{0.70}{\scriptstyle\pm 0.07}$ $\mathbf{0.16}{\scriptstyle\pm 0.01}$ $\mathbf{0.28}{\scriptstyle\pm 0.01}$ $\mathbf{0.46}{\scriptstyle\pm 0.03}$ $\underline{0.56}{\scriptstyle\pm 0.06}$ $\underline{0.50}{\scriptstyle\pm 0.02}$
EventFlow ($\alpha=0$) $0.69{\scriptstyle\pm 0.17}$ $2.01{\scriptstyle\pm 0.40}$ $0.26{\scriptstyle\pm 0.01}$ $0.47{\scriptstyle\pm 0.03}$ $1.23{\scriptstyle\pm 0.07}$ $0.66{\scriptstyle\pm 0.03}$ $0.80{\scriptstyle\pm 0.05}$
E.2 MMDs
In Tables 17 and 18, we report the MMD values appearing in the unconditional experiment (i.e., Table 3 in the main paper) with standard deviations. These are omitted from the main paper for the sake of space.
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
Data $1.3$ $0.6$ $0.4$ $3.1$ $2.6$ $3.6$ $3.1$
Hawkes1 Hawkes2 NSP NSR SC SR
Data $1.3$ $1.3$ $1.8$ $3.0$ $5.7$ $1.1$
E.3 MSEs
In Table 19, we evaluate the performance of our model when forecasting only a single subsequent event. That is, given a history $\mathcal{H}$, we evaluate the MSE between the first true event time following this history and the first event time generated by each model conditioned on $\mathcal{H}$. The results are reported in Table 19. Generally, all of the methods show similar results on this metric, despite there being clear differences between methods on the multi-step task. We believe this serves to further highlight the necessity of moving beyond single-step prediction tasks.
Table 19: MSE values evaluating one-step-ahead forecasting performance. Mean values $\pm$ one standard deviation are reported over five random seeds. The lowest MSE on each dataset is indicated and bold, and the second lowest is indicated by an underline.
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
IFTPP $0.85{\scriptstyle\pm 0.05}$ $\underline{0.32}{\scriptstyle\pm 0.03}$ $0.0047{\scriptstyle\pm 0.0006}$ $0.22{\scriptstyle\pm 0.03}$ $1.74{\scriptstyle\pm 0.10}$ $1.24{\scriptstyle\pm 0.16}$ $1.11{\scriptstyle\pm 0.17}$
NHP $0.89{\scriptstyle\pm 0.09}$ $0.53{\scriptstyle\pm 0.24}$ $\mathbf{0.0022}{\scriptstyle\pm 0.0007}$ $0.31{\scriptstyle\pm 0.12}$ $2.00{\scriptstyle\pm 0.30}$ $1.30{\scriptstyle\pm 0.26}$ $\underline{1.03}{\scriptstyle\pm 0.35}$
Diffusion $\mathbf{0.61}{\scriptstyle\pm 0.10}$ $0.33{\scriptstyle\pm 0.04}$ $\underline{0.0037}{\scriptstyle\pm 0.0012}$ $0.23{\scriptstyle\pm 0.14}$ $\mathbf{1.30}{\scriptstyle\pm 0.21}$ $\mathbf{0.86}{\scriptstyle\pm 0.18}$ $\mathbf{0.92}{\scriptstyle\pm 0.20}$
Add-and-Thin $0.86{\scriptstyle\pm 0.05}$ $\mathbf{0.30}{\scriptstyle\pm 0.04}$ $0.0043{\scriptstyle\pm 0.0007}$ $\underline{0.21}{\scriptstyle\pm 0.03}$ $\underline{1.53}{\scriptstyle\pm 0.14}$ $1.16{\scriptstyle\pm 0.16}$ $1.20{\scriptstyle\pm 0.14}$
EventFlow (25 NFEs) $\underline{0.68}{\scriptstyle\pm 0.02}$ $1.06{\scriptstyle\pm 0.09}$ $0.028{\scriptstyle\pm 0.0015}$ $\mathbf{0.19}{\scriptstyle\pm 0.01}$ $2.30{\scriptstyle\pm 0.22}$ $\underline{0.90}{\scriptstyle\pm 0.05}$ $1.30{\scriptstyle\pm 0.03}$
EventFlow (10 NFEs) ${0.60}{\scriptstyle\pm 0.02}$ $1.01{\scriptstyle\pm 0.31}$ $0.0113{\scriptstyle\pm 0.0016}$ ${0.16}{\scriptstyle\pm 0.01}$ $1.93{\scriptstyle\pm 0.07}$ ${0.77}{\scriptstyle\pm 0.05}$ ${1.14}{\scriptstyle\pm 0.07}$
EventFlow (1 NFE) $0.65{\scriptstyle\pm 0.16}$ $1.30{\scriptstyle\pm 0.61}$ $0.0445{\scriptstyle\pm 0.0104}$ $0.18{\scriptstyle\pm 0.04}$ $1.52{\scriptstyle\pm 0.17}$ $0.95{\scriptstyle\pm 0.17}$ $1.03{\scriptstyle\pm 0.24}$
E.4 Forecasting TPPs
In Table 20, we report the sequence distance values appearing in the forecasting experiment (i.e., Figure 2 and Table 1 in the main paper) with standard deviations. We additionally report an ablation where we use the true value of $n$, rather than sampling $n\sim p_{\phi}(n\mid\mathcal{H})$. This serves to measure how much room for improvement remains from the event count predictor. In general, we see that this benchmark is still not saturated, with further gains being possible especially on datasets with long sequences (e.g., the Reddit-C and Reddit-S datasets).
Table 20: Sequence distance (9) between the forecasted and ground-truth event sequences on a held-out test set. Lower is better. We report the mean $\pm$ one standard deviation over five random seeds. The lowest mean distance on each dataset is indicated in bold, and the second lowest by an underline.
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
IFTPP $4.2{\scriptstyle\pm 0.7}$ $25.6{\scriptstyle\pm 2.3}$ $61.2{\scriptstyle\pm 3.2}$ $5.1{\scriptstyle\pm 0.4}$ $2.9{\scriptstyle\pm 0.2}$ $2.1{\scriptstyle\pm 0.2}$ $3.4{\scriptstyle\pm 0.2}$
NHP $2.8{\scriptstyle\pm 0.1}$ $31.0{\scriptstyle\pm 1.4}$ $95.7{\scriptstyle\pm 0.7}$ $4.5{\scriptstyle\pm 0.3}$ $3.4{\scriptstyle\pm 0.5}$ $\underline{1.8}{\scriptstyle\pm 0.1}$ $\underline{3.0}{\scriptstyle\pm 0.2}$
Diffusion $5.4{\scriptstyle\pm 1.2}$ $25.7{\scriptstyle\pm 0.9}$ $80.3{\scriptstyle\pm 11.4}$ $4.6{\scriptstyle\pm 0.7}$ $\underline{2.4}{\scriptstyle\pm 0.2}$ $\underline{1.8}{\scriptstyle\pm 0.1}$ $3.3{\scriptstyle\pm 0.7}$
Add-and-Thin $\underline{2.5}{\scriptstyle\pm 0.04}$ $\underline{22.2}{\scriptstyle\pm 4.6}$ $\underline{34.3}{\scriptstyle\pm 0.4}$ $\underline{3.7}{\scriptstyle\pm 0.1}$ $3.1{\scriptstyle\pm 0.2}$ $\underline{1.8}{\scriptstyle\pm 0.1}$ $\underline{3.0}{\scriptstyle\pm 0.2}$
EventFlow (25 NFEs) $\mathbf{2.0}{\scriptstyle\pm 0.03}$ $\mathbf{15.8}{\scriptstyle\pm 2.7}$ $\mathbf{16.0}{\scriptstyle\pm 0.2}$ $\mathbf{3.2}{\scriptstyle\pm 0.1}$ $\mathbf{1.4}{\scriptstyle\pm 0.01}$ $\mathbf{1.3}{\scriptstyle\pm 0.02}$ $\mathbf{1.9}{\scriptstyle\pm 0.1}$
EventFlow (10 NFEs) $2.0{\scriptstyle\pm 0.03}$ $15.8{\scriptstyle\pm 2.7}$ $15.8{\scriptstyle\pm 0.2}$ $3.1{\scriptstyle\pm 0.1}$ $1.3{\scriptstyle\pm 0.02}$ $1.3{\scriptstyle\pm 0.1}$ $1.9{\scriptstyle\pm 0.1}$
EventFlow (1 NFE) $2.0{\scriptstyle\pm 0.01}$ $15.8{\scriptstyle\pm 2.7}$ $15.8{\scriptstyle\pm 0.2}$ $3.2{\scriptstyle\pm 0.2}$ $1.4{\scriptstyle\pm 0.03}$ ${1.8\scriptstyle\pm 0.3}$ $1.9{\scriptstyle\pm 0.1}$
EventFlow (25 NFEs, true $n$) $1.2{\scriptstyle\pm 0.01}$ $5.5{\scriptstyle\pm 0.3}$ $8.8{\scriptstyle\pm 0.2}$ $1.8{\scriptstyle\pm 0.02}$ $0.7{\scriptstyle\pm 0.01}$ $0.7{\scriptstyle\pm 0.02}$ $1.1{\scriptstyle\pm 0.02}$